if 4/7 of a quantity is 24 then 5/6 of the same quantity is
A. 10 B. 14 C. 13 D. 4 E. 5



Answer by palanisamy(496)

Given, 4/7 of a quantity is 24
Then 5/6 of the same quantity is = (5/6)*24/(4/7)
= (5/6)*24*(7/4)
= 35
Another Method
Let the quantity be x.
Given, (4/7)x = 24
x = 24*(7/4)
x = 6*7
x = 42
Therefore (5/6)x = (5/6)*42 = 5*7 = 35